OpenStudy (anonymous):

Solve: 8x2 + 64x = 0

OpenStudy (ali2x2):

http://www.tiger-algebra.com/drill/8x2_64x=0/

OpenStudy (midhun.madhu1987):

8x2 + 64x = 0 Take 8x common factor: 8x(x + 8)= 0 which gives 8x = 0 or x+8 = 0 so what are the 2 values of x?